OpenStudy (anonymous):

Arrange these numbers from least to greatest: -5/9, -7/9, -4/5, -3/7 A. -7/9, -5/8, -3/7, -4/5 B. -4/5, -7/9, -5/8, -3/7 C. -4/5, -5/8, -7/9, -3/7 D. -3/7, -4/5, -5/8, -7/9

OpenStudy (anonymous):

ok so -3/7 = -.42 -4/5= -.8 -5/8= .-625 -7/9= -.77

OpenStudy (anonymous):

-3/7, -5/8, -7/9, -4/5 but they are all negative sooooo its the opposite way around

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thanks!